OpenStudy (helpmeplease14):

How do circumference and area change when the radius of a circle is doubled? Provide an example.

OpenStudy (helpmeplease14):

@KendrickLamar2014

OpenStudy (helpmeplease14):

@rational

OpenStudy (helpmeplease14):

@TheSmartOne

OpenStudy (helpmeplease14):

@kidrah69

OpenStudy (campbell_st):

do a little calculation.. find the area and circumference is 1 unit... then double the radius r = 2 then calculate the circumference and area...compare the results.

OpenStudy (helpmeplease14):

?

OpenStudy (campbell_st):

use the formulae for 1. circumference \[C = 2 \pi r\] 2, Area \[A = \pi r^2\] firstly with a radius of r = 1 unit find the area and circumference of the circle then use a radius of r = 2 units to find the area and circumference of the circle then compare you answers fo r = 1 and r = 2 the information for the comparison of the circumference will be different to the information for the area. hope it helps

OpenStudy (helpmeplease14):

OK thx

OpenStudy (helpmeplease14):

Forget the question mark

OpenStudy (anonymous):

Lol Shaniqua! Guess who it is?